# Computing Large 2D/3D Convex Regions of Obstacle-Free Space for Robotic Planning

### v1

@@ III-A Problem Formulation
Consider that a convex robot is navigating in an obstacle-rich $n$-dimensional environment, where $n={2,3}$. The geometrical shape of the robot is given by the $\mathcal{V}$-representation [33] of a convex polytope, i.e., convex combination of a finite number of points
@@
in which $v_{1},\dots,v_{s}\in\mathbb{R}^{n}$ are allowed to be redundant, which means that they are not required to only contain the extreme points of $\mathcal{Q}$. The obstacle region $\mathcal{O}$ is, albeit nonconvex, assumed to be the union of convex obstacles $\mathcal{O}=\cup_{i=1}^{R}\mathcal{O}_{i}$, of which the $i$-th one is determined by $s_{i}$ points
@@
We require no collision between the robot $\mathcal{Q}$ and the obstacle region $\mathcal{O}$, thus implying $\mathcal{Q}\cap\mathcal{O}=\varnothing$.
Our problem is to compute an obstacle-free convex polytope $\mathcal{P}$ which is required to contain the robot $\mathcal{Q}$ while excluding all obstacles $\mathcal{O}_{i}$. Besides, $\mathcal{P}$ should have the largest possible volume within a prescribed region of interest. For convenience, we define that the boundaries of the prescribed region are considered as obstacles and are encoded into $\mathcal{O}$, which makes the volume of obstacle-free space surrounding $\mathcal{Q}$ bounded, as shown in Fig. 1. Concluding above requirements yields the optimization
$\max{\operatorname*{vol}({\mathcal{P}})},~{}\mathrm{s.t.}~{}\mathcal{Q}\subseteq\mathcal{P},~{}\mathcal{O}\subseteq\operatorname*{coint}({\mathcal{P}}),$ (3)
where $\operatorname*{vol}({\mathcal{P}})$ denotes the volume of the convex polytope $\mathcal{P}$ and $\operatorname*{coint}({\cdot})$ denotes the complement of the interior of a set. Note that the solution set of (3) will never be empty since the robot $\mathcal{Q}$ itself is already a feasible solution.
Figure 1: Illustration of the computation process of FIRI, corresponding to the Restrictive Inflation and MVIE calculation steps depicted in Algorithm. 1.
Input: $\mathcal{Q},\mathcal{O},\mathcal{E}_{0},\rho$
Output: $\mathcal{P}$
@@
2       $k\leftarrow 0$ repeat
3             $k\leftarrow k+1$ /* Restrictive Inflation starts */ $\mathcal{R}\leftarrow\{{1,\dots,R}\}$ $\mathcal{E}\leftarrow\mathcal{E}_{k-1},~{}\bar{\mathcal{P}}\leftarrow\mathbb{R}^{n}$ $\bar{\mathcal{Q}}\leftarrow D_{\mathcal{E}}^{-1}A_{\mathcal{E}}^{\mathrm{T}}({\mathcal{Q}-b_{\mathcal{E}}})$ foreach $i\in\mathcal{R}$ do
4                   $\bar{\mathcal{O}}_{i}\leftarrow D_{\mathcal{E}}^{-1}A_{\mathcal{E}}^{\mathrm{T}}({\mathcal{O}_{i}-b_{\mathcal{E}}})$ $a_{i}\leftarrow\operatorname*{arg\,max}_{a\in\mathbb{R}^{n}}{a^{\mathrm{T}}a},$ ${}~{}~{}~{}~{}~{}~{}\mathrm{s.t.}~{}\bar{\mathcal{Q}}\subseteq\mathcal{H}({a}),~{}\bar{\mathcal{O}}_{i}\subseteq\operatorname*{coint}({\mathcal{H}({a})})$
@@ 2       $k\leftarrow 0$ repeat
7                   $j\leftarrow\operatorname*{arg\,min}_{i\in\mathcal{R}}{a_{i}^{\mathrm{T}}a_{i}}$ $\bar{\mathcal{P}}\leftarrow\bar{\mathcal{P}}\cap\mathcal{H}({a_{j}})$$\mathcal{R}\leftarrow\mathcal{R}\backslash\left\{{i\in\mathcal{R}~{}\big{|}~{}\bar{\mathcal{O}}_{i}\subseteq\operatorname*{coint}({\mathcal{H}({a_{j}})})}\right\}$
8            until $\mathcal{R}=\varnothing$$\mathcal{P}_{k}\leftarrow A_{\mathcal{E}}D_{\mathcal{E}}\bar{\mathcal{P}}+b_{\mathcal{E}}$ /* Restrictive Inflation ends */ /* MVIE starts */ $\mathcal{E}_{k}\leftarrow\mathcal{E}^{*}({\mathcal{P}_{k}})$ /* MVIE ends */
9      until $\operatorname*{vol}({\mathcal{E}_{k}})\leq(1+\rho)\operatorname*{vol}({\mathcal{E}_{k-1}})$ return $\mathcal{P}_{k}$
@@
III-B Proposed Algorithm
For the optimization (3), dealing with constraints on nonconvex sets directly is intractable [34] and no known polynomial complexity algorithm has been given so far for the case of $n>3$. Thus we only seek a high-quality solution to (3) instead of the global optimal solution. Additionally, we optimize a reasonable lower bound of vol$(\mathcal{P})$, which is straightforward to evaluate, to maximize the original objective function vol$(\mathcal{P})$ whose computation is NP-hard [17]. As adopted in [1], we also choose the the volume of the Maximum Volume Inscribed Ellipsoid (MVIE) of $\mathcal{P}$ for this lower bound. An ellipsoid is defined by
=$\mathcal{E}=\left\{{p~{}\big{|}~{}p=A_{\mathcal{E}}D_{\mathcal{E}}x+b_{\mathcal{E}},~{}\|{x}\|=1}\right\},$ (4)
where $A_{\mathcal{E}}\in\mathbb{R}^{n\times n}$ is orthonormal, $D_{\mathcal{E}}\in\mathbb{R}^{n\times n}$ is diagonal and non-negative, and $b_{\mathcal{E}}\in\mathbb{R}^{n}$. The diagonal elements of $D_{\mathcal{E}}$ correspond to the lengths of the semi-axes of $\mathcal{E}$.
The proposed novel algorithm Fast Iterative Region Inflation (FIRI), as shown in Algorithm 1. FIRI takes the robot $\mathcal{Q}$, the obstacles $\mathcal{O}$, an initial ellipsoid $\mathcal{E}_{0}$, and a precision parameter $\rho\in(0,1)$ as inputs. Any ellipsoid that is strictly contained in $\mathcal{Q}$ can be used to initialize the algorithm. FIRI iteratively executes two modules Restrictive Inflation and sollving MVIE within its outer loop (line 1-1). The former takes an ellipsoid as input and expands the ellipsoid to obtain a new convex hull, and the latter takes the convex hull as input and computes its MVIE.
In Restrictive Inflation (line 1-1), For each obstacle $\mathcal{O}_{i}$, we maximally inflate the input ellipsoid $\mathcal{E}$ under the constraint that there exists a halfspace which does not contain the obstacle yet includes both the inflated ellipsoid and robot $\mathcal{Q}$. Then we use these halfspaces to forms a new convex polytope as the output of Restrictive Inflation. Specifically, $\mathcal{Q}$ and $\mathcal{O}_{i}$ are transformed (line 1 and 1) by the inverse affine map determined by the ellipsoid $\mathcal{E}$ generated in the previous iteration. Since $\mathcal{Q}$ and $\mathcal{O}_{i}$ are both in $\mathcal{V}$-representation, their images after transformation are still the convex combinations of the images of their vertices, i.e.,
@@ III-B Proposed Algorithm
The affine transformation can be massively parallelized for all vertices. It is important that in the transformed space the ellipsoid $\mathcal{E}$ is transformed to a unit ball $\mathcal{B}$ as
=$\mathcal{B}=\left\{{p~{}|~{}\|{p}\|=1}\right\},$ (7)
whose collision check is far cheaper to handle than the ellipsoid $\mathcal{E}$. Then the maximum inflation ratio of the unit ball $\mathcal{B}$ for each transformed obstacle $\bar{\mathcal{O}}_{i}$ is computed in the first inner loop (line 1-1). Compared to the methods that lake the managibility to include the robot [1, 4], the proposed Restrictive Inflation compute a halfspace for each obstacle $\bar{\mathcal{O}}_{i}$ such that it contains $\bar{\mathcal{Q}}$ and the corresponding inflated ball $\mathcal{B}_{i}$ yet excludes $\bar{\mathcal{O}}_{i}$ while maximizing the $\mathcal{B}_{i}$ which is defined by
$\mathcal{B}_{i}=\left\{{p~{}|~{}\|{p}\|=a_{i}^{\mathrm{T}}a_{i}}\right\},$ (8)
where $a_{i}$ is the only contact point of the halfspace with the inflated ball $\mathcal{B}_{i}$. The halfspace $\mathcal{H}({a_{i}})$ is defined as
$\mathcal{H}({a_{i}}):=\left\{{x\in\mathbb{R}^{n}~{}\big{|}~{}a_{i}^{\mathrm{T}}x\leq a_{i}^{\mathrm{T}}a_{i}}\right\}.$ (9)
Then Restrictive Inflation can be written as an optimization as shown in line 1. The second inner loop (line 1-1) iteratively finds the closest allowable halfspace, adds it into $\bar{\mathcal{P}}$, and removes unused obstacles. If all obstacles in $\mathcal{I}$ have been processed, a feasible polytope in $\mathcal{H}$-representation [33] can be formed in the transformed space. Finally, a new polytope $\mathcal{P}_{k}$ is generated by recovering $\bar{\mathcal{P}}$ to the original space (line 1). Thus, the output of FIRI is an obstacle-free convex polytope in $\mathcal{H}$-representation, i.e., intersection of $m$ halfspaces
@@
where $A_{\mathcal{P}}\in\mathbb{R}^{m\times n}$ and $b_{\mathcal{P}}\in\mathbb{R}^{m}$.
For computing the MVIE of a convex polytope (line 1), as long as a bounded $\mathcal{P}$ has nonempty interior, its unique MVIE [35] is determined by
*$\mathcal{E}^{*}({\mathcal{P}})=\max_{A_{\mathcal{E}},D_{\mathcal{E}},b_{\mathcal{E}}}{\operatorname*{vol}({\mathcal{E}})},~{}\mathrm{s.t.}~{}\mathcal{E}\subseteq\mathcal{P}.$ (11)
FIRI continuously optimizes the lower bound of the objective function of the original problem (3). During the iterative computation, this algorithm always ensures the feasibility of the solution, which means the generated convex polytope $\mathcal{P}_{k}$ maintains satisfacting the constraints of the original problem as $k$ increases. Moreover, it maintains the monotonicity of the volume of ellipsoid $\mathcal{E}_{k}$ as the lower bound of the volume of the convex polytope $\mathcal{P}_{k}$.
For the feasibility, each halfspace solved in Restrictive Inflation (line 1) satisfies the original constraints, thus the new convex polytope $\mathcal{P}_{k}$ formed by the intersection of these halfspaces is guaranteed to be feasible, given by
$\mathcal{Q}\subseteq\mathcal{P}_{k},~{}\mathcal{O}\subseteq\operatorname*{coint}({\mathcal{P}_{k}}).$ (12)
Additionally, as defined in Sec III-A, the prescribed region of interest is bounded and the boundaries of this region are encoded into $\mathcal{O}$, thus the feasibility also indicates that the generated convex polytope $\mathcal{P}_{k}$ is closed.
For the monotonicity, in the $k$-th iteration, we inflate the unit ball $\mathcal{B}$ for each transformed obstacle $\bar{\mathcal{O}}_{i}$ in Restrictive Inflation, which always ensures the contact point satisfies $\|{a_{i}}\|\geq 1$. Thus, in the transformed space, the MVIE of $\bar{\mathcal{P}}$ acquired from the second inner loop (line 1-1) can never be smaller than the unit ball $\mathcal{B}$ as
$\operatorname*{vol}({\mathcal{E}^{*}({\bar{\mathcal{P}}})})\geq\operatorname*{vol}({\mathcal{B}}).$ (13)
Since the relative size relationship between the MVIE of $\bar{\mathcal{P}}$ and the unit ball $\mathcal{B}$ is invariant to affine maps, and $\mathcal{B}$ is obtained by applying the inverse affine map to the input ellipsoid $\mathcal{E}_{k-1}$, in the original space (line 1 and 1) we have
$\operatorname*{vol}({\mathcal{E}_{k}})\geq\operatorname*{vol}({\mathcal{E}_{k-1}}).$ (14)

### v2

@@ III-A Problem Formulation
Consider a convex seed in an obstacle-rich $n$-dimensional environment, where $n\in\{2,3\}$. The geometrical shape of the seed is given by the $\mathcal{V}$-representation [38] of a convex polytope, i.e., convex combination of a finite number of points
@@
in which $v_{1},\dots,v_{s}\in\mathbb{R}^{n}$ are allowed to be redundant, which means that they are not required to only contain the extreme points of $\mathcal{Q}$. The obstacle region $\mathcal{O}$ is, albeit nonconvex, assumed to be the union of convex obstacles $\mathcal{O}=\cup_{i=1}^{N}\mathcal{O}_{i}$, of which the $i$-th one is determined by $s_{i}$ points
@@
We require no collision between the seed $\mathcal{Q}$ and the obstacle region $\mathcal{O}$, thus implying $\mathcal{Q}\cap\mathcal{O}=\varnothing$.
Our problem is to compute an obstacle-free convex polytope $\mathcal{P}$ which is required to contain the seed $\mathcal{Q}$ while excluding all obstacles $\mathcal{O}$. Besides, $\mathcal{P}$ should have the largest possible volume within a prescribed region of interest. For convenience, we define that the boundaries of the prescribed region are considered as obstacles and are encoded into $\mathcal{O}$, which makes the volume of obstacle-free space surrounding $\mathcal{Q}$ bounded. Concluding above requirements yields the optimization
$\max_{\mathcal{P}}~{}{\operatorname*{vol}({\mathcal{P}})},~{}~{}\mathrm{s.t.}~{}\mathcal{Q}\subseteq\mathcal{P},~{}\mathcal{O}\subseteq\operatorname*{coint}({\mathcal{P}}),$ (3)
where $\operatorname*{vol}({\mathcal{P}})$ denotes the volume of the convex polytope $\mathcal{P}$ and $\operatorname*{coint}({\cdot})$ denotes the complement of the interior of a set. Note that the solution set of (3) will never be empty since the seed $\mathcal{Q}$ itself is already a feasible solution.
[…]
III-B Algorithm Overview
Even if we simplify the optimization (3) by disregarding the constraint of including the seed, specifically $\mathcal{Q}\subseteq\mathcal{P}$, it will be transformed into a generalization of the challenging potato peeling problem, which has an $O(N^{7})$ approach [39] in 2-D case and is NP-hard [40] in 3-D case. Not to mention that bringing in this constraint complicates the optimization (3) even more than potato peeling. To achieve a balance between efficiency and quality, we only seek a high-quality feasible solution to (3) efficiently instead of the global optimal solution. Additionally, since the computation of the objective function vol$(\mathcal{P})$ of (3) is at least as hard as an NP-complete problem [41], we optimize a reasonable lower bound of vol$(\mathcal{P})$, which is straightforward to evaluate, to maximize the original objective function. As adopted in [1], we also choose the the volume of the MVIE of $\mathcal{P}$ as this lower bound.
To efficiently maximize the lower bound of the volume of the convex polytope $\mathcal{P}$ while satisfying the constraints of the original problem (3), we propose a novel algorithm called Fast Iterative Region Inflation (FIRI) as shown in Algorithm 1. Any ellipsoid that is strictly contained in $\mathcal{Q}$ can be used to initialize the algorithm. FIRI iteratively executes two modules: RsI and MVIE within its outer loop (Line 1-1). In the $k$-th iteration, the former module takes an ellipsoid as input and expands the ellipsoid to obtain a new convex polytope $\mathcal{P}_{k}$, and the latter module takes the convex polytope $\mathcal{P}_{k}$ as input and computes its MVIE $\mathcal{E}_{k}$. As illustrated in Fig. 2, the output of these two modules serves as the input for each other. Then we provide the details of these two modules and the convergence and manageability of FIRI.
In this paper, we define an ellipsoid $\mathcal{E}$ by
=$\mathcal{E}=\left\{{p~{}\big{|}~{}p=A_{\mathcal{E}}D_{\mathcal{E}}x+b_{\mathcal{E}},~{}x\in\mathbb{R}^{n},~{}\|{x}\|=1}\right\},$ (4)
where $A_{\mathcal{E}}\in\mathbb{R}^{n\times n}$ is orthonormal, $D_{\mathcal{E}}\in\mathbb{R}^{n\times n}$ is diagonal and positive-definite, and $b_{\mathcal{E}}\in\mathbb{R}^{n}$. The diagonal elements of $D_{\mathcal{E}}$ correspond to the lengths of the semi-axis of $\mathcal{E}$.
Notion:  number of obstacles $N$,
           parameters of ellipsoid $\mathcal{E}$$A_{\mathcal{E}},D_{\mathcal{E}},b_{\mathcal{E}}$,
           halpspace $\mathcal{H}({a})$ defined in (9)
Input:   seed $\mathcal{Q}$, obstacles $\mathcal{O}$, initial ellipsoid $\mathcal{E}_{0}$,
           threshold $\rho$
Output: convex polytope $\mathcal{P}$
@@ III-B Algorithm Overview
2       $k\leftarrow 0$
3       repeat
4             $k\leftarrow k+1$
5             /* RsI starts */
6             $\mathcal{I}\leftarrow\{{1,\dots,N}\}$
7             $\mathcal{E}\leftarrow\mathcal{E}_{k-1},~{}\bar{\mathcal{P}}\leftarrow\mathbb{R}^{n}$
8             $\bar{\mathcal{Q}}\leftarrow D_{\mathcal{E}}^{-1}A_{\mathcal{E}}^{\mathrm{T}}({\mathcal{Q}-b_{\mathcal{E}}})$
9             foreach $i\in\mathcal{I}$ do
10                   $\bar{\mathcal{O}}_{i}\leftarrow D_{\mathcal{E}}^{-1}A_{\mathcal{E}}^{\mathrm{T}}({\mathcal{O}_{i}-b_{\mathcal{E}}})$
11                   $a_{i}\leftarrow\operatorname*{arg\,max}_{a\in\mathbb{R}^{n}}{a^{\mathrm{T}}a},$ ${}~{}~{}~{}~{}~{}~{}\mathrm{s.t.}~{}\bar{\mathcal{Q}}\subseteq\mathcal{H}({a}),~{}\bar{\mathcal{O}}_{i}\subseteq\operatorname*{coint}({\mathcal{H}({a})})$
12
@@ 7             $\mathcal{E}\leftarrow\mathcal{E}_{k-1},~{}\bar{\mathcal{P}}\leftarrow\mathbb{R}^{n}$
15                   $j\leftarrow\operatorname*{arg\,min}_{i\in\mathcal{I}}{a_{i}^{\mathrm{T}}a_{i}}$
16                   $\bar{\mathcal{P}}\leftarrow\bar{\mathcal{P}}\cap\mathcal{H}({a_{j}})$
17                   \$\mathcal{I}\leftarrow\mathcal{I}\backslash\left\{{i\in\mathcal{I}~{}\big{|}~{}\bar{\mathcal{O}}_{i}\subseteq\operatorname*{coint}({\mathcal{H}({a_{j}})})}\right\}$
18
19            until $\mathcal{I}=\varnothing$
20            $\mathcal{P}_{k}\leftarrow A_{\mathcal{E}}D_{\mathcal{E}}\bar{\mathcal{P}}+b_{\mathcal{E}}$
21             /* RsI ends */
22             /* MVIE starts */
23             $\mathcal{E}_{k}\leftarrow\mathcal{E}^{*}({\mathcal{P}_{k}})$
24             /* MVIE ends */
25
26      until $\operatorname*{vol}({\mathcal{E}_{k}})\leq(1+\rho)\operatorname*{vol}({\mathcal{E}_{k-1}})$
27       return $\mathcal{P}_{k}$
@@ 23             $\mathcal{E}_{k}\leftarrow\mathcal{E}^{*}({\mathcal{P}_{k}})$
III-B1 RsI
For each obstacle $\mathcal{O}_{i}$, we maximally inflate the input ellipsoid $\mathcal{E}$ under the constraint that there exists a halfspace which does not contain the obstacle yet includes both the inflated ellipsoid and seed $\mathcal{Q}$. Then we use these halfspaces to forms a new polytope as the output of RsI.
Specifically, $\mathcal{Q}$ and $\mathcal{O}_{i}$ are transformed (Line 1 and 1) by the inverse affine map determined by the ellipsoid $\mathcal{E}$ generated in the previous iteration. Since $\mathcal{Q}$ and $\mathcal{O}_{i}$ are both in $\mathcal{V}$-representation, their images after transformation are still the convex combinations of the images of their vertices, i.e.,
@@ III-B1 RsI
It is important that in the transformed space the ellipsoid $\mathcal{E}$ is transformed to a unit ball $\mathcal{B}$ as
=$\mathcal{B}=\left\{{x\in\mathbb{R}^{n}~{}|~{}\|{x}\|=1}\right\},$ (7)
whose collision check is far cheaper to handle than ellipsoids.
Then we calculate a restrictive halfspace for each transformed obstacle $\bar{\mathcal{O}}_{i}$ in the first inner loop (Line 1-1). Compared to the methods that lack the manageability to include the seed [1, 2], the proposed RsI requires that the halfspace contains $\bar{\mathcal{Q}}$ and the corresponding inflated ball $\mathcal{B}_{i}$ yet excludes $\bar{\mathcal{O}}_{i}$. Additionally, it aims to maximize the size of $\mathcal{B}_{i}$ which is defined by
$\mathcal{B}_{i}=\left\{{x\in\mathbb{R}^{n}~{}|~{}\|{x}\|=a_{i}^{\mathrm{T}}a_{i}}\right\},$ (8)
where $a_{i}$ is the only contact point of the boundary of the halfspace with the inflated ball $\mathcal{B}_{i}$. Then the halfspace $\mathcal{H}({a_{i}})$ can be defined as
$\mathcal{H}({a_{i}})=\left\{{x\in\mathbb{R}^{n}~{}\big{|}~{}a_{i}^{\mathrm{T}}x\leq a_{i}^{\mathrm{T}}a_{i}}\right\}.$ (9)
The calculation of the restrictive halfspace can be written as an optimization as shown in Line 1, for which we provide an efficient solution method in Sec. IV. In the left diagram of Fig. 2(a), we provide a specific example of the restrictive halfspace computation in the transformed space.
In the second inner loop (Line 1-1), we generate a new convex polytope based on the obtained halfspaces. We iteratively finds the closest halfspace $\mathcal{H}({a_{i}})$, adds it into $\bar{\mathcal{P}}$, and then remove the halfspaces corresponding to the obstacles outside of $\mathcal{H}({a_{i}})$. Until all halfspaces are processed, an obstacle-free polytope in $\mathcal{H}$-representation [38] can be formed in the transformed space. Then a new polytope $\mathcal{P}_{k}$ is obtained by recovering $\bar{\mathcal{P}}$ to the original space (Line 1). Thus, the output of FIRI is an obstacle-free convex polytope in $\mathcal{H}$-representation, i.e., intersection of $m$ halfspaces
@@
[…]
III-B2 MVIE
As long as a closed $\mathcal{P}$ has nonempty interior, its unique MVIE [42] $\mathcal{E}^{*}({\mathcal{P}})$ can be determined by solving
$\max_{A_{\mathcal{E}},D_{\mathcal{E}},b_{\mathcal{E}}}{\operatorname*{vol}({\mathcal{E}})},~{}\mathrm{s.t.}~{}\mathcal{E}\subseteq\mathcal{P},$ (11)
which is employed in Line 1. To solve MVIE, we propose efficient methods in Sec. V and Sec. VI so that the computational overhead of MVIE will no longer be a stumbling block that prevents this monotonically inflating MVIE approach from being applied to real-time scenarios [1].
III-B3 Manageability and Convergence of FIRI
During the iterative computation, FIRI always ensures the feasibility of the solution, which means the generated convex polytope $\mathcal{P}_{k}$ maintains satisfacting the constraints of the original problem as $k$ increases. Moreover, it maintains the monotonicity of the volume of MVIE $\mathcal{E}_{k}$, which is the lower bound of the volume of the convex polytope $\mathcal{P}_{k}$. Then by analyzing the feasibility and monotonicity of the output of FIRI, we explain the manageability and convergence they bring about.
For the feasibility, since each halfspace computed in RsI (Line 1) satisfies the original constraints, the new convex polytope $\mathcal{P}_{k}$ formed by the intersection of these halfspaces is guaranteed to be feasible :
$\mathcal{Q}\subseteq\mathcal{P}_{k},~{}\mathcal{O}\subseteq\operatorname*{coint}({\mathcal{P}_{k}}),$ (12)
which gives the manageability of FIRI. Additionally, as defined in Sec III-A, the prescribed region of interest is bounded, whose boundaries are encoded into $\mathcal{O}$, thus the feasibility also indicates that RsI always generates a closed polytope $\mathcal{P}_{k}$.
For the monotonicity, in the $k$-th iteration, we inflate the unit ball $\mathcal{B}$ for each transformed obstacle $\bar{\mathcal{O}}_{i}$ in RsI, which always ensures the contact point $a_{i}$ satisfies $\|{a_{i}}\|\geq 1$. Thus, in the transformed space, $\bar{\mathcal{P}}$ acquired from the second inner loop (Line 1-1) always holds $\mathcal{B}\subseteq\bar{\mathcal{P}}$, which means that in the original space (Line 1) we have $\mathcal{E}_{k-1}\subseteq\mathcal{P}_{k}$. As $\mathcal{E}_{k}$ is the largest ellipsoid contained by $\mathcal{P}_{k}$, we have
$\operatorname*{vol}({\mathcal{E}_{k-1}})\leq\operatorname*{vol}({\mathcal{E}_{k}}).$ (13)